Solve the given initial value problem. y - 12 +55y - 114y = 0; y(0) = 1, y(0) = 0, y (0) = 0 y(t) = 0
Answer #1

SOLUTION : =) D=d dt y–12 +55Y –114 y = 0 (D3 - 12D+55 D – 114) y=0 Pts auxiliary equation m3_ 12m2 + ssm-114=0 by hit aY (t) = 36c eft + Cosuiot (92+ 3 VTo C₂ + 3NT0 C3-1062) + sindłot ( 963 - 3√ To C₂ - 3 VTo G-10) y (t) = 369 €64+ 3+ COSNTo